O Week Four Quiz Question 23 Not yet answered Points out of s.00 Flag question The heights of 18-year-old men are approximately normally distributed with mean 68 inches and standard deviation 3 inches. What is the probability that the average height of a sample of twenty 18-year-old men will be less than 69 inches? Round your answer to four decimal places Select one: O a. 0.8064 o b. 09032 О С. 0.4516 O d. 0.4032 O e.o.0968 Next page Previous page O Type here to search 8 7
Answer #1
for normal distribution z score =(X-μ)/σ
here mean=       μ= 68
std deviation   =σ= 3.0000
sample size       =n= 20
std error=σ=σ/√n= 0.6708
probability = P(X<69) = P(Z<1.49)= 0.9320

option B is most appropriate